the perimeter of a college basketball court is 96m and the lenght is 14m more than the width. what are the dimensions.
Answer by mananth(16946)

let width =x
length = x+14
..
Perimeter = 96 = 2*(L+W)
96=2*(X+X+14)
96=2*(2x+14)
96=4x+28
96-28=4x
68=4x
/4
68/4=x
17=x the width
length = 17+14= 31m
...
CHECK
17+17+31+31=96